# Learning Stochastic Bridges for Video Object Removal via Video-to-Video Translation

### v2

@@ 2 Related Works
Recently, bridge models (Chen et al., 2021; Tong et al., 2024; Liu et al., 2023; Zhou et al., 2023a; Chen et al., 2023; Zheng et al., 2024; He et al., 2024; De Bortoli et al., 2021; Peluchetti, 2023) have garnered significant attention for their ability to transcend the limitations of the Gaussian prior inherent in standard diffusion models. By establishing a direct stochastic path between two arbitrary distributions, bridge models have demonstrated the clear superiority of a data-to-data generation paradigm over traditional noise-to-data approaches, particularly in tasks such as image-to-image translation (Liu et al., 2023; Zhou et al., 2023a) and speech synthesis (Chen et al., 2023; Li et al., 2025a). Notably, FrameBridge (Wang et al., 2025) successfully extended this methodology to image-to-video (I2V) synthesis by introducing SNR-aligned fine-tuning to stabilize the learning process.
Figure 1: The framework of BridgeRemoval. The video inputs are projected into latent space using a frozen VAE. Unlike standard diffusion, we employ a VP-SDE Bridge formulation to interpolate a trajectory ($z_{t}$) directly from the source video prior ($z_{\text{src}}$) to the clean target ($z_{\text{tgt}}$). The DiT-based model is conditioned on text embeddings ($c_{\text{text}}$) and a spatial input formed by concatenating the mask ($z_{\text{M}}$) with the source latent ($z_{\text{src}}$), optimized via a velocity-matching objective ($\mathcal{L}_{\text{bridge}}$).
@@ 3.1 Brownian Bridge
The standard Brownian Bridge (Albergo et al., 2023; Li et al., 2023) serves as a foundational instance of such processes, conditioned to start at a specific point $x_{1}$ and end at $x_{0}$. Unlike standard diffusion which maps data to Gaussian noise, a Brownian Bridge interpolates between two data points. Assuming a constant diffusion coefficient, the intermediate state $X_{t}$ follows a simple linear interpolation perturbed by noise:
@@ 4.1 Latent Encoding and Dual-Role Conditioning
Given a training triplet consisting of a source video $V_{\text{src}}$, a target video $V_{\text{tgt}}$, and a binary mask $M$, we first project the visual data into a compressed latent space using a pre-trained VAE encoder $\mathcal{E}$:
@@
Simultaneously, the text prompt $P$ is encoded into embeddings $c_{\text{text}}$ via a T5 encoder.
@@ 4.4 Adaptive Mask Modulation
While the bridge formulation effectively anchors the generation trajectory to the source video, thereby ensuring superior background consistency, it introduces a critical trade-off. The source video acts as a strong prior, providing structural guidance that is beneficial for stability. However, this rigid adherence to the source distribution can become detrimental when removing large objects.
To mitigate this limitation, we propose adaptive mask modulation (AMM). This mechanism is designed to dynamically recalibrate the feature influence based on the mask, enhancing the utilization of valid background context while relaxing the constraints on the inpainting region.
Let $h\in\mathbb{R}^{N\times D}$ denote the input patch embeddings to the Transformer, and $M\in\{0,1\}^{N\times 1}$ be the corresponding binary mask (where $1$ denotes the object region and 0 the background). We introduce a learnable modulation parameter $\gamma$. The modulated embeddings $h^{\prime}$ are computed as:
$h^{\prime}=h\odot\left(M+(1-M)\cdot(1+\gamma)\right).$ (15)
This operation differentially scales the input features:
  • Object Region ($M=1$): Features remain unchanged ($h^{\prime}=h$), preserving the necessary noise information for the bridge process.
  • Background Region ($M=0$): Features are amplified by a factor of $(1+\gamma)$.
By adaptively boosting the magnitude of embeddings in the unmasked background regions, AMM effectively counterbalances the strong source prior within the masked regions, providing the necessary generative flexibility to reconstruct large occlusions while maintaining strict fidelity to the surrounding environment.
@@ 5.1 Experimental Setup
Training Data. We construct a training dataset totaling approximately 47,000 video pairs from two complementary sources. First, we utilize the ROSE dataset (Miao et al., 2025), which consists of 16,000 synthetically generated samples. By leveraging 3D engines to toggle object visibility, ROSE provides perfectly aligned triplets of source videos, target backgrounds, and masks. While this offers high-quality supervision, our statistical analysis indicates a distribution bias towards static, rigid objects, limiting the model’s generalization to complex motion. To address this limitation, we introduce a supplementary real-world composite dataset comprising 31,000 video pairs. This subset is explicitly designed to improve robustness on dynamic agents, containing 15,000 human-centric videos and 16,000 videos of common dynamic objects (e.g., animals, vehicles). The data pipeline adopts a composition-based approach: we employ SAM2 (Ravi et al., 2024) to extract high-fidelity foregrounds from real-world videos and composite them onto clean backgrounds. A Qwen3-VL (Bai et al., 2025b) based filtering mechanism is then applied to ensure visual coherence and discard low-quality samples.
@@
For all 150 videos, object masks are generated using SAM2 (Ravi et al., 2024). Furthermore, to support methods such as VACE (Jiang et al., 2025) that require a textual description of the target video (object removed), we employ Qwen3-VL (Bai et al., 2025b) to generate a detailed caption for each video (object removed), and performed the same process on the DAVIS dataset. In summary, the DAVIS dataset and BridgeRemoval-Bench are used to evaluate the performance of various algorithms.

### v3

@@ 2 Related Works
Recently, bridge models (Chen et al., 2021; Tong et al., 2024; Liu et al., 2023; Zhou et al., 2023a; Chen et al., 2023; Zheng et al., 2025; He et al., 2024; De Bortoli et al., 2021; Peluchetti, 2023) have garnered significant attention for their ability to transcend the limitations of the Gaussian prior inherent in standard diffusion models. By establishing a direct stochastic path between two arbitrary distributions, bridge models have demonstrated the clear superiority of a data-to-data generation paradigm over traditional noise-to-data approaches, particularly in tasks such as image-to-image translation (Liu et al., 2023; Zhou et al., 2023a) and speech synthesis (Chen et al., 2023; Li et al., 2025a). Notably, FrameBridge (Wang et al., 2025) successfully extended this methodology to image-to-video (I2V) synthesis by introducing SNR-aligned fine-tuning to stabilize the learning process.
Figure 1: The framework of BridgeRemoval. The video inputs are projected into latent space using a frozen VAE encoder. Unlike standard diffusion, we employ a VP-SDE Bridge formulation to interpolate a trajectory ($z_{t}$) directly from the source video prior ($z_{\text{src}}$) to the clean target ($z_{\text{tgt}}$). The DiT-based model is conditioned on text embeddings ($c_{\text{text}}$) and a spatial input formed by concatenating the mask ($z_{\text{M}}$) with the source latent ($z_{\text{src}}$), optimized via a velocity-matching objective ($\mathcal{L}_{\text{bridge}}$).
@@ 3.1 Brownian Bridge
The standard Brownian Bridge (Albergo et al., 2025; Li et al., 2023) serves as a foundational instance of such processes, conditioned to start at a specific point $x_{1}$ and end at $x_{0}$. Unlike standard diffusion which maps data to Gaussian noise, a Brownian Bridge interpolates between two data points. Assuming a constant diffusion coefficient, the intermediate state $X_{t}$ follows a simple linear interpolation perturbed by noise:
@@ 4.1 Latent Encoding and Dual-Role Conditioning
Given a training triplet consisting of a source video $V_{\text{src}}$, a target video $V_{\text{tgt}}$, and a binary mask $M$, we first project the visual data into a compressed latent space using a pre-trained VAE encoder $\mathcal{E}$ (Wan et al., 2025):
@@
Simultaneously, the text prompt $P$ is encoded into embeddings $c_{\text{text}}$ via a T5 encoder (Raffel et al., 2020).
@@ 4.4 Adaptive Mask Modulation
While the bridge formulation effectively anchors the generation trajectory to the source video, this rigid adherence to the source distribution can become detrimental when removing large objects. To mitigate this limitation, we propose Adaptive Mask Modulation (AMM), a spatially-aware feature modulation mechanism. The mask $M\in\mathbb{R}^{C_{m}\times F\times H\times W}$ is then projected into a latent feature space through a learnable patch embedding:
$F_{M}=\phi_{M}(M)\in\mathbb{R}^{D\times F^{\prime}\times H^{\prime}\times W^{\prime}}$ (15)
where $\phi_{M}$ is a 3D convolutional layer. From the mask features $F_{M}$, we predict spatially-varying modulation parameters through two $1\times 1\times 1$ convolutional projections:
$\gamma=f_{\gamma}(F_{M}),\quad\beta=f_{\beta}(F_{M})$ (16)
where $\gamma,\beta\in\mathbb{R}^{D\times F^{\prime}\times H^{\prime}\times W^{\prime}}$ represent the scale and shift parameters, respectively.
Let $h\in\mathbb{R}^{D\times F^{\prime}\times H^{\prime}\times W^{\prime}}$ denote the input patch embeddings after the initial projection. The modulated embeddings $h^{\prime}$ are computed as:
$h^{\prime}=h\odot(1+\gamma)+\beta$ (17)
Scale modulation $(1+\gamma)$ adaptively amplifies or attenuates features based on their spatial relationship to the mask, enabling the model to learn where to preserve original information. Shift modulation $(\beta)$ injects position-aware guidance into the feature, providing explicit spatial cues about the inpainting region boundaries and structure.
By incorporating rich spatial mask information through learnable modulation, AMM enables the model to adaptively adjust its behavior across different spatial regions, effectively counterbalancing the strong source prior within masked regions while maintaining strict fidelity to the surrounding environment.
@@ 5.1 Experimental Setup
Training Data. We construct a training dataset totaling approximately 47,000 video pairs from two complementary sources. First, we utilize the ROSE dataset (Miao et al., 2025), which consists of 16,000 synthetically generated samples. By leveraging 3D engines to toggle object visibility, ROSE provides perfectly aligned triplets of source videos, target backgrounds, and masks. While this offers high-quality supervision, our statistical analysis indicates a distribution bias towards static, rigid objects, limiting the model’s generalization to complex motion. To address this limitation, we introduce a supplementary real-world composite dataset comprising 31,000 video pairs. This subset is explicitly designed to improve robustness on dynamic agents, containing 15,000 human-centric videos and 16,000 videos of common dynamic objects (e.g., animals, vehicles). The data pipeline adopts a composition-based approach: we employ SAM2 (Ravi et al., 2025) to extract high-fidelity foregrounds from real-world videos and composite them onto clean backgrounds. A Qwen3-VL (Bai et al., 2025b) based filtering mechanism is then applied to ensure visual coherence and discard low-quality samples.
@@
For all 150 videos, object masks are generated using SAM2 (Ravi et al., 2025). Furthermore, to support methods such as VACE (Jiang et al., 2025) that require a textual description of the target video (object removed), we employ Qwen3-VL (Bai et al., 2025b) to generate a detailed caption for each video (object removed), and performed the same process on the DAVIS dataset. In summary, the DAVIS dataset and BridgeRemoval-Bench are used to evaluate the performance of various algorithms.